# Convergence of Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex functions

### v3

@@ 4 Convergence Properties of Decentralized Stochastic Subgradient-based Methods
As discussed in Section 2.3, the class of definable functions is wide enough to enclose the loss functions of common nonsmooth neural networks. Moreover, [36] demonstrates that there exists a definable conservative field for any definable loss function constructed by definable blocks, within which the results yielded by Automatic Differentiation (AD) algorithms are contained. Henceforth, Assumption 4.1(1) is mild and reasonable in practical scenarios.
For conciseness, we denote $D_{F_{i,l}}:=D_{F_{i}(\cdot;{\bm{s}}_{l})}$. According to [36, Corollary 4] and Assumption 4.1(1),
@@ 4.1 Decentralized stochastic subgradient descent
In this subsection, we establish the global convergence for decentralized stochastic subgradient descent (DSGD) by showing that it fits into our proposed framework (DSM). The detailed algorithm of DSGD is presented in Algorithm 1. Intuitively, for each $i\in[d]$, the consensus term $\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}$ enforces ${\bm{x}}_{i,k}$ towards the mean over all the agents (i.e., $\frac{1}{d}\sum_{i\in[d]}{\bm{x}}_{i,k}$). Moreover, as ${\bm{d}}_{i,k}$ refers to the (mini-batch) stochastic subgradient of $f_{i}$ at ${\bm{x}}_{i,k}$, the term $-\eta_{k}{\bm{d}}_{i,k}$ can be viewed as a single SGD step to minimize $f_{i}$ in the $i$-th agent.
@@
Let the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. We set ${\bm{Z}}_{k}:={\bm{X}}_{k}$, ${\bm{H}}_{k}:=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}:={\bm{D}}_{k}-{\bm{H}}_{k}$. To demonstrate that Algorithm 1 fits into our framework (DSM), it suffices to check the update scheme (4.3) is a special case of (DSM) satisfying Assumption 3.1, and verify the validity of Assumption 3.8 or 3.12 with different sampling techniques. Assumption 3.1(1) follows immediately from Assumption 4.2(2) or 4.3(2); Assumption 3.1(2) follows quickly from local boundedness of the conservative field $D_{F_{i,l}}$. Nevertheless, how to verify (1.5) varies across different cases. For case (1) in Theorem 4.4, we set $\Phi_{i}:=D_{f_{i}}$. In this case, the nonnegative sequence $\{\delta_{k}\}$ in (1.5) is not subject to any other restrictions, hence we may select a sufficiently large one to meet the practical requirement for ${\bm{H}}_{k}$ and $\Phi_{i}$. This choice makes Assumption (1.5) hold trivially. For case (2), we need to carefully choose a diminishing $\{\delta_{k}\}$ so that it conforms to the conditions in Theorem 3.15.
@@
Furthermore, according to [36, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, combining Assumption 4.1(1) and Proposition 2.7 yields $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
[…]
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [43] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables.
In this subsection, motivated by the generalized SGD with momentum analyzed in [43], we introduce the decentralized generalized stochastic subgradient descent method with momentum (DGSGD-M) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [43], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
@@
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Hence, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, when we choose $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$, Algorithm 2 corresponds to DSGD with momentum; when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$, which can be viewed as a decentralized variant of SignSGD analyzed in [51], and denoted as DSignSGD. Interested readers could refer to [43] for more details on the choices of the auxiliary function $\phi$.
@@
Now we check the validity of (1.5) and Assumption 3.1, and then show Assumption 3.8 and 3.12 are satisfied for case (1) and (2), respectively. From the local boundedness of $D_{f_{i}}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$, if $\{{\bm{X}}_{k}\}$ is bounded. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. This illustrates that Assumption 3.1(2) holds.
@@
Furthermore, from the definition of $\Phi_{i}$ and (1.5), we attain that $\Phi({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau D_{f}({\bm{x}})\end{smallmatrix}\right]$. Then together with [43, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion $\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\Phi({\bm{x}},{\bm{y}})$, and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Leveraging the definability of $f$ and $D_{f}$, it follows that the set $\{\psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.1 holds for Algorithm 2.
We set $\mathcal{U}_{i,j}({\bm{x}},{\bm{y}}):=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\frac{1}{|\mathcal{B}_{i,j}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,j}}D_{F_{i,l}}({\bm{x}})\end{smallmatrix}\right]$, where $\mathcal{B}_{i,j}$ for $10\leq j\leq N-$ 1 are subsets of equal size, and set $\rho_{k}:=\eta_{k}\|{\bm{T}}_{k}\|+2\sum_{j=0}^{k-1}\lambda_{2}^{k-1-j}\eta_{j}\|{\bm{T}}_{j}\|$. Then (4.4) is a special form of (DSM) with (3.8). Moreover, Assumption 4.2 implies that Assumption 3.8 is satisfied for case (1). With the choice of $\xi_{i,k}:=\{\textit{Draw out }\mathcal{B}_{i,k}\textit{ from agent }i\}$, and
@@
(4.4) is a particular form of (DSM) with (3.18), and Assumption 3.12 is satisfied for case (2). Plugging Lemma 3.6 and Theorem 3.14, we can prove that $\delta_{k}\to 0$, as $k\to\infty$ with arbitrary high probability  [^2] Notice that Theorem 3.14 does not require $\delta_{k}$ to diminish.. Employing Theorem 3.11 and 3.15, we can derive the results in Theorem 4.7.
Theorem 4.8 illustrates how Algorithm 2 avoids spurious critical points, whose proof is presented in Appendix A for simplicity.
@@ 4.3 Decentralized SGD with gradient-tracking technique
In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM) with (3.8) or (3.18). Therefore, based on our theoretical results exhibited in Theorem 3.11 and 3.15, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
In the rest of this subsection, we denote $,{\bm{V}}_{k}=[{\bm{v}}_{1,k},$$\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. We condense the assertion that DSGD-T can fit in our framework (DSM) with different noise settings, along with convergence analysis, into a single theorem.
@@
[…]
@@
Assumption 3.1(1) follows quickly from Assumption 4.2 or 4.3. (1.5) holds vacuously with sufficiently large $\delta_{k}$ and $\Phi_{i}:=D_{f_{i}}$ for case (1). Moreover for case (2), followed by
$\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k}),$ (4.6)
it is easy to deduce that
@@
Furthermore, according to [36, Section 6], we can assert that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $D_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we have verified the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
Following the same construction of $\mathcal{U}_{i,k}({\bm{x}})$, $\rho_{k}$, $\xi_{i,k}$ and $\chi_{i}$ as in the proof of Theorem 4.4, we could verify Assumption 3.8 and 3.12 hold for case (1) and case (2), respectively. Henceforth, the update scheme (4.5) is a variant of (DSM) with either (3.8) or (3.18). Utilizing our theoretical results in Theorem 3.11 and 3.15, we completes the proof.

### v4

@@ 4 Convergence Properties of Decentralized Stochastic Subgradient-based Methods
As discussed in Section 2.3, the class of definable functions is wide enough to enclose the loss functions of common nonsmooth neural networks. Moreover, [bolte2021conservative] demonstrates that there exists a definable conservative field for any definable loss function constructed by definable blocks, within which the results yielded by Automatic Differentiation (AD) algorithms are contained. Henceforth, Assumption 4.1(1) is mild and reasonable in practical scenarios.
For conciseness, we denote $D_{F_{i,l}}:=D_{F_{i}(\cdot;{\bm{s}}_{l})}$. According to [bolte2021conservative, Corollary 4] and Assumption 4.1(1),
@@ 4.1 Decentralized stochastic subgradient descent
In this subsection, we establish the global convergence for decentralized stochastic subgradient descent (DSGD) by showing that it fits into our proposed framework (DSM). The detailed algorithm of DSGD is presented in Algorithm 1. Intuitively, for each $i\in[d]$, the consensus term $\sum_{j\in\mathcal{N}_{i}}{\bm{W}}(i,j){\bm{x}}_{j,k}$ enforces ${\bm{x}}_{i,k}$ towards the mean over all the agents (i.e., $\frac{1}{d}\sum_{i\in[d]}{\bm{x}}_{i,k}$). Moreover, as ${\bm{d}}_{i,k}$ refers to the (mini-batch) stochastic subgradient of $f_{i}$ at ${\bm{x}}_{i,k}$, the term $\eta_{k}{\bm{d}}_{i,k}$ can be viewed as a single SGD step to minimize $f_{i}$ in the $i$-th agent.
@@
Let the filtration $\mathcal{F}_{k}:=\sigma(\{{\bm{X}}_{l}:l\leq k\})$. We set ${\bm{Z}}_{k}:={\bm{X}}_{k}$, ${\bm{H}}_{k}:=\mathbb{E}[{\bm{D}}_{k}|\mathcal{F}_{k}]$ and $\Xi_{k+1}:={\bm{D}}_{k}-{\bm{H}}_{k}$. To demonstrate that Algorithm 1 fits into our framework (DSM), it suffices to check the update scheme (4.3) is a special case of (DSM) satisfying Assumption 3.1, and verify the validity of Assumption 3.9 or 3.13 with different sampling techniques. Assumption 3.1(1) follows immediately from Assumption 4.2(2) or 4.3(2); Assumption 3.1(2) follows quickly from local boundedness of the conservative field $D_{F_{i,l}}$. Nevertheless, how to verify (1.5) varies across different cases. For case (1) in Theorem 4.4, we set $\Phi_{i}:=D_{f_{i}}$. In this case, the nonnegative sequence $\{\delta_{k}\}$ in (1.5) is not subject to any other restrictions, hence we may select a sufficiently large one to meet the practical requirement for ${\bm{H}}_{k}$ and $\Phi_{i}$. This choice makes Assumption (1.5) hold trivially. For case (2), we need to carefully choose a diminishing $\{\delta_{k}\}$ so that it conforms to the conditions in Theorem 3.16.
@@
Furthermore, according to [bolte2021conservative, Section 6] and the path-differentiability of $f$, it follows that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, combining Assumption 4.1(1) and Proposition 2.7 yields $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we verify the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
With the choice of $\mathcal{U}_{i,j}({\bm{x}}):=\frac{1}{|\mathcal{B}_{i,j}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,j}}D_{F_{i,l}}({\bm{x}})$, where $\mathcal{B}_{i,j},0\leq j\leq N-1$ are subsets of equal size, and $\rho_{k}:=0$, Assumption 3.9 is satisfied for case (1); With the choice of $\xi_{i,k}:=\{\textit{Draw out }\mathcal{B}_{i,k}\textit{ from }$  $\textit{ agent }i\}$, $\delta_{k}:=0$ and $\chi_{i}({\bm{x}},\xi_{i,k})\in\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}})-D_{f_{i}}({\bm{x}})$, Assumption 3.13 is satisfied for case (2). Finally, setting $\mathcal{X}_{0}$ as a compact set which contains ${\bm{x}}_{0}$, and leveraging Theorem 3.12 and 3.16, we completes our proof. ∎
@@ 4.2 Decentralized generalized SGD with momentum
Momentum is a technique to accelerate the subgradient descent method in the relevant direction and dampen oscillations. In the single-agent setting, [xiao2023convergence] investigates the global convergence properties of generalized SGD with momentum, where an auxiliary function $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is introduced to normalize the update directions of the primal variables.
In this subsection, motivated by the generalized SGD with momentum analyzed in [xiao2023convergence], we introduce the decentralized generalized stochastic subgradient descent method with momentum (DGSGD-M) and present the details in Algorithm 2. Different from DSGD in Algorithm 1, Algorithm 2 introduces the auxiliary variables $\{{\bm{y}}_{i,k}:k\in\mathbb{N}\}$ to track the first-order moment of the stochastic subgradients of $f_{i}$. Therefore, Algorithm 2 requires two rounds of communication to perform local consensus to $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$ and $\{{\bm{y}}_{1,k},\ldots,{\bm{y}}_{d,k}\}$, respectively. Moreover, similar to the generalized SGD analyzed in [xiao2023convergence], Algorithm 2 introduces an auxiliary function $\phi$ to normalize the update directions of its primal variables $\{{\bm{x}}_{1,k},\ldots,{\bm{x}}_{d,k}\}$.
@@
The auxiliary function $\phi$ in Algorithm 2 determines how the update directions of the variables $\{{\bm{x}}_{k}\}$ are regularized. Hence, Algorithm 2 yields different variants of DSGD-type methods with different choices of the auxiliary function $\phi$. For example, when we choose $\phi({\bm{y}})=\frac{1}{2}\left\|{\bm{y}}\right\|^{2}$, Algorithm 2 corresponds to DSGD with momentum; when we select $\phi({\bm{y}})=\left\|{\bm{y}}\right\|_{1}$, Algorithm 2 employs the sign map to regularize the update directions of the variables $\{{\bm{x}}_{k}\}$, which can be viewed as a decentralized variant of SignSGD analyzed in [bernstein2018signsgd], and denoted as DSignSGD. Interested readers could refer to [xiao2023convergence] for more details on the choices of the auxiliary function $\phi$.
Furthermore, recognizing signSGD as a limiting variant of Adam [chen2023symbolic], we extend our analysis in Appendix C to demonstrate how Adam-type methods can be adapted to our framework, hence establishing convergence guarantees for these methods under the nonsmooth nonconvex settings.
@@
Now we check the validity of (1.5) and Assumption 3.1, and then show Assumption 3.9 and 3.13 are satisfied for case (1) and (2), respectively. From the local boundedness of $D_{f_{i}}$, we can conclude that $\{{\bm{D}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. Together with the update scheme of ${\bm{Y}}_{k}$, we can conclude that $\left\|{\bm{Y}}_{k+1}\right\|\leq(1-\tau\eta_{k})\left\|{\bm{Y}}_{k}\right\|+\tau\eta_{k}\left\|{\bm{D}}_{k+1}\right\|$, hence there exists a constant $M_{Y}>0$ such that $\sup_{k\geq 0}\left\|{\bm{Y}}_{k}\right\|\leq\max\{\left\|{\bm{Y}}_{0}\right\|,\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\}<M_{Y}$, if $\{{\bm{X}}_{k}\}$ is bounded. Then combining with local boundedness of $\partial\phi$, it can be referred that $\{{\bm{T}}_{k}\}$ is uniformly bounded almost surely whenever $\{{\bm{X}}_{k}\}$ is bounded. This illustrates that Assumption 3.1(2) holds.
@@
Furthermore, from the definition of $\Phi_{i}$ and (1.5), we attain that $\Phi({\bm{x}},{\bm{y}})=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau D_{f}({\bm{x}})\end{smallmatrix}\right]$. Then together with [xiao2023convergence, Proposition 4.5] and the path-differentiability of $f$, we can conclude that $\psi({\bm{x}},{\bm{y}})=f({\bm{x}})+\frac{1}{\tau}\phi({\bm{y}})$ is the Lyapunov function of the differential inclusion $\left(\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t},\frac{\mathrm{d}{\bm{y}}}{\mathrm{d}t}\right)\in-\Phi({\bm{x}},{\bm{y}})$, and admits $\{({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ as its stable set. Leveraging the definability of $f$ and $D_{f}$, it follows that the set $\{\psi({\bm{x}},{\bm{y}})\in\mathbb{R}^{n}\times\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}}),{\bm{y}}={\bm{0}}\}$ is finite in $\mathbb{R}$. Therefore, Assumption 3.1 holds for Algorithm 2.
We set $\mathcal{U}_{i,j}({\bm{x}},{\bm{y}}):=\left[\begin{smallmatrix}\partial\phi({\bm{y}})\\ \tau{\bm{y}}-\tau\frac{1}{|\mathcal{B}_{i,j}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,j}}D_{F_{i,l}}({\bm{x}})\end{smallmatrix}\right]$, where $\mathcal{B}_{i,j}$ for $0\leq j\leq N-1$ are subsets of equal size, and set $\rho_{k}:=\eta_{k}\|{\bm{T}}_{k}\|+2\sum_{j=0}^{k-1}\lambda_{2}^{k-1-j}\eta_{j}\|{\bm{T}}_{j}\|$. Then (4.4) is a special form of (DSM) with (3.8). Moreover, Assumption 4.2 implies that Assumption 3.9 is satisfied for case (1). With the choice of $\xi_{i,k}:=\{\textit{Draw out }\mathcal{B}_{i,k}\textit{ from agent }i\}$, and
@@
(4.4) is a particular form of (DSM) with (3.18), and Assumption 3.13 is satisfied for case (2). Plugging Lemma 3.7 and Theorem 3.15, we can prove that $\delta_{k}\to 0$, as $k\to\infty$ with arbitrary high probability  [^2] Notice that Theorem 3.15 does not require $\delta_{k}$ to diminish.. Employing Theorem 3.12 and 3.16, we can derive the results in Theorem 4.7.
Theorem 4.8 illustrates how Algorithm 2 avoids spurious critical points, whose proof is presented in Appendix A for simplicity. We also provide a numerical example in Appendix B to demonstrate that DSGD-m can avoid spurious critical points.
@@ 4.3 Decentralized SGD with gradient-tracking technique
In this subsection, we prove that DSGD-T can be enclosed by our proposed framework (DSM) with (3.8) or (3.18). Therefore, based on our theoretical results exhibited in Theorem 3.12 and 3.16, we can establish the global convergence of DSGD-T in the minimization of nonsmooth path-differentiable functions.
@@
In the rest of this subsection, we denote ${\bm{V}}_{k}=[{\bm{v}}_{1,k},$ $\ldots,{\bm{v}}_{d,k}]$ at $k$-th iteration in Algorithm 3. We condense the assertion that DSGD-T can fit in our framework (DSM) with different noise settings, along with convergence analysis, into a single theorem.
@@
$\leq{}$ $3\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|+\left(\sup_{k\geq 0}\left\|{\bm{D}}_{k}\right\|\right)\cdot\sum_{i=1}^{k}\left\|({\bm{W}}-{\bm{I}}_{d}){\bm{W}}^{k+1-i}\right\|$
$\leq{}$ $3M_{D}+\sum_{i=1}^{\infty}2M_{D}\alpha^{i}\leq\left(3+\frac{2}{1-\alpha}\right)M_{D}.$
@@
Assumption 3.1(1) follows quickly from Assumption 4.2 or 4.3. (1.5) holds vacuously with sufficiently large $\delta_{k}$ and $\Phi_{i}:=D_{f_{i}}$ for case (1). Moreover for case (2), followed by $\mathbb{E}[{\bm{d}}_{i,k}|\mathcal{F}_{k}]\in\mathbb{E}_{\mathcal{B}_{i,k}}[\frac{1}{|\mathcal{B}_{i,k}|}\sum_{{\bm{s}}_{l}\in\mathcal{B}_{i,k}}D_{F_{i,l}}({\bm{x}}_{i,k})]=D_{f_{i}}({\bm{x}}_{i,k})$, it is easy to deduce that
@@
Furthermore, according to [bolte2021conservative, Section 6], we can assert that $f$ is the Lyapunov function of the differential inclusion $\frac{\mathrm{d}{\bm{x}}}{\mathrm{d}t}\in-D_{f}({\bm{x}})$, and admits the stable set $\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$. In addition, the definability of $f$ and $D_{f}$ illustrates that $\{f({\bm{x}}):{\bm{0}}\in D_{f}({\bm{x}})\}$ is finite in $\mathbb{R}$. As a result, we have verified the validity of Assumption 3.1 with $\Phi:=D_{f}$, $\psi:=f$, and $\mathcal{A}:=\{{\bm{x}}\in\mathbb{R}^{n}:{\bm{0}}\in D_{f}({\bm{x}})\}$.
Following the same construction of $\mathcal{U}_{i,k}({\bm{x}})$, $\rho_{k}$, $\xi_{i,k}$ and $\chi_{i}$ as in the proof of Theorem 4.4, we could verify Assumption 3.9 and 3.13 hold for case (1) and case (2), respectively. Henceforth, the update scheme (4.5) is a variant of (DSM) with either (3.8) or (3.18). Utilizing our theoretical results in Theorem 3.12 and 3.16, we completes the proof.